OpenStudy (ittybitty14):

what is 3 square root of 135? A. 5 3 root of 2 B. 2 3 root of 5 C. 5 3 root of 3 D. 3 3 root of 5

OpenStudy (ittybitty14):

MEDAL AND FAN!!!!!!!!!!!

OpenStudy (welshfella):

first find the prime factors of 135

OpenStudy (ittybitty14):

I think it's 3^3 x 5

OpenStudy (welshfella):

its 3*3*3*5 can we get a perfect square from these numbers?

OpenStudy (ittybitty14):

Yes?

OpenStudy (welshfella):

oh i misread the question! you want the cube root of 135 (\[\sqrt[3]{135}\]

OpenStudy (mathmale):

If the index of your root is 2, you have a square root. If the index is 3, you have a cube root (not a 3 square root).

OpenStudy (welshfella):

= \[\sqrt[3]{27} * \sqrt[3]{5}\]

OpenStudy (mathmale):

I agree with welshfella: you want the third root of 135. Look at the way he has factored 135. What do you think was his purpose in doing that?

OpenStudy (welshfella):

yes the '3 square root' confused me.

OpenStudy (mathmale):

What is the "nice" third root of 27? Does 5 have a "nice" third root (i. e., with NO decimal fraction)?